The sum of 3 consecutive even integers is 294 what are the integers?
Found 2 solutions by ewatrrr, Alan3354:
Answer by ewatrrr(24785)

 
Hi
3 consecutive even integers
Let x, (x+2) and (x+4) represent the 3 consecutive even integers
Question states***
x + (x+2) + (x+4) = 294
3x = 288
x = 96 The 3 consecutive even integers are 96,98,100
CHECKING our Answer***
96+98+100 = 294

Answer by Alan3354(69443)

294/3 = 98, the middle integer
--> 96, 98, 100